A bag contains five red balls, three white balls and two green balls. If three balls are selected at random, with replacement, find the probability that the first two are red and third one is green.
Answer by stanbon(57407)

A bag contains five red balls, three white balls and two green balls. If three balls are selected at random, with replacement, find the probability that the first two are red and third one is green.
----------------
1st red = 5/10
2nd red = 4/9
----------------
3rd green = 2/8
================
P(rrg) = (5/10)(4/9)(2/8) = (4/9)(1/8) = 1/18
=================
Cheers,
Stan H.